OpenStudy (anonymous):

Connections Academy algebra 2A part 1?

OpenStudy (anonymous):

can you post the actual question?

OpenStudy (anonymous):

1. which system of equations is shown by the graph below? What is the solution? a.-x+5y=2 x-2y=2 solution:(-2,0) b.-x+5y=2 x-2y=-2 solution:(-2,0 c.x+5y=2 x+2y=2 solution:(0,-2) d.x+5y=-2 x+2y+-2 solution:(0,-2)